OpenStudy (anonymous):

7. Find S8 for the geometric series 256 + 64 + 16 + 4 +… a. 357.62 b. 341. 33 c. 338. 25 d. 324.13

OpenStudy (amistre64):

theres a nice little formula, but you need to know the common ratio to use it

OpenStudy (amistre64):

if i recall it cirrectly, it goes\[\frac{1-r^n}{1-r}\]

OpenStudy (amistre64):

the brute math method is to list all 8 terms and add them up, but youd stil need to determine the common ratio

OpenStudy (anonymous):

r = 1/4

OpenStudy (anonymous):

use that and find

OpenStudy (amistre64):

how does r = 1/4 ?

OpenStudy (anonymous):

WHAT AMIRSTRE SAID I MEAN WHYYYYYYYYY WOULD YOU DO THAT YAHOO WHY

OpenStudy (anonymous):

since it is in gp

OpenStudy (anonymous):

r = 64/256

OpenStudy (anonymous):

=1/4

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

in general, each new term is the product of the one before it and some common value (r) \[a_{n+1}=a_n*r\] to find r, we divide both dies by an \[\frac{a_{n+1}}{a_n}=r\]and that determines the common ratio

OpenStudy (anonymous):

in gp each term is multiplied........

OpenStudy (amistre64):

take any 2 of the numbers that are in a row, 16 and 4 and stack them last over first: 4/16 = r = 1/4

OpenStudy (amistre64):

to test this r out, lets dbl chk: 256*1/4 = 64 64*1/4 = 16 16*1/4 = 4 4*1/4 = 1 1*1/4 = 1/4 1/4 * 1/4 = 1/16 etc ....

OpenStudy (amistre64):

the rest is to plug everything into the formula for summation:\[\frac{1-r^n}{1-r}\to\ \frac{1-(1/4)^8}{1-1/4}=\frac{something}{3/4}\]

OpenStudy (anonymous):

It is a decimal though which is not one of the answers....

OpenStudy (amistre64):

fractions are decimals, you just have to simplify it

OpenStudy (amistre64):

a calculator helps at times

OpenStudy (amistre64):

1/4 = .25 might be useful

OpenStudy (amistre64):

we might also want to make it easier, add up the inter parts and just use summation onthe fractions remaining

OpenStudy (anonymous):

i think there is a tiny mistake here

OpenStudy (amistre64):

256*1/4 = 64 64*1/4 = 16 16*1/4 = 4 4*1/4 = 1 .................................. 85 n=4 left

OpenStudy (anonymous):

where??

OpenStudy (amistre64):

i forgot a 256 along the way eh

OpenStudy (anonymous):

yeah there

OpenStudy (amistre64):

256 85 ---- 341 + n=3 left

OpenStudy (amistre64):

341+ something smaller than 1 would seem to be b right?

OpenStudy (anonymous):

I'm taking that! thank you for extensively helping me lol

OpenStudy (amistre64):

lol, youre welcome ;) good luck

OpenStudy (anonymous):

yes it is definitely 341 and a bit

OpenStudy (anonymous):

\[\sum_{k=0}^7256(\frac{1}{4})^n=256\times \frac{1-(\frac{1}{4})^8}{\frac{3}{4}}\]